# Semantic Web technologies in sensor-based personal health monitoring systems: A systematic mapping study

### v2

@@ 5.1.5 The role of Semantic Web technologies
Semantic Web technologies are critical in the achievement of semantic interoperability among the selected systems. Many of the systems make use of terminologies such as SNOMED CT and ICD through ontologies, and access knowledge graphs such as DrugBank and Diseasome which are published as linked data. This allows the systems to access and reason with standardised health domain knowledge. Additionally, many systems use ontologies and knowledge graphs to represent sensor data and health records, which can enhance process interoperability. Semantic Web technologies can also contribute to syntactic interoperability, albeit in an indirect capacity. For instance, El-Sappagh et al. [84] map FHIR resources to an ontology, allowing for interoperability between their proposed system and hospital information systems that use FHIR. However, Semantic Web technologies do not inherently provide support for technical interoperability since they operate at a higher, more abstract level to formally represent and derive meaning from the data. Therefore, in order to achieve technical interoperability, health monitoring systems must leverage data transmission protocols and devices.
@@ 5.2 Context awareness
An important aspect of health monitoring is the ability to take context into consideration, which is critical for situation analysis since contextual information enhances sensor data and supports its interpretation. Consider a case where an individual’s heart rate is suddenly elevated. If the individual is engaged in exercise, the increased heart rate is expected. However, if the individual is at rest, this could be a cause for alarm. Therefore, health monitoring systems must be able to adapt based on the context of the individual being monitored. The four most common aspects of context are location, time, identity (of a person or agent), and activity (or events) [122, 123].
@@ 5.2.1 Location
Ye et al. [122] highlight three types of locations that can be represented: symbolic locations, coordinate locations, and regions. The systems proposed by Akhtar et al. [76], Chiang and Liang [82], and Vadillo et al. [108] keep track of the different rooms in a house where an individual may be, while those proposed by Khozouie et al. [96], Titi et al. [107], and Zeshan et al. [111] indicate more generally the place the monitored individual is (for example, “home” or “hospital”). These are symbolic locations. One purpose of such locations is to allow the systems to suggest relevant services based on the type of space currently occupied, as is the case in the system proposed by Chiang and Liang [82]. In the system proposed by Hristoskova et al. [92], the clinician’s location (i.e. the room they occupy in a hospital) is used to determine which device to send notifications to, optimising for the closest device. Likewise, Zeshan et al. [111] determine the closeness between the monitored individual and their caregivers in order to select which caregiver or clinician to notify in case of abnormal sensor observations. This is similar to the system proposed by Alti et al. [80], which supports a GPS sensor that captures the current coordinates of the monitored individual. In this system, location is used to select devices closest to the user from which to deploy health services so as to increase efficiency and minimise inter-device communication costs. Coordinate locations also serve the purpose of alerting caregivers and emergency services of the exact location of a person in the event of a medical emergency, as is suggested by Rhayem et al. [105]. The system proposed by Hadjadj and Halimi [89] integrates health monitoring in the public transport system, and therefore includes location sensors in public transportation vehicles. The final type of location is regions, which are geometrical two- or three-dimensional representations of locations [122]. This type of location is used in the system proposed by Kim et al. [97] in order to advise users of region-specific situations, such as adverse or dangerous weather. Similarly, El-Sappagh et al. [84] use the spatial region class from the Basic Formal Ontology to represent the patient’s current location, as well as the placement of the sensors. Despite the importance of location as an aspect of context, less than half of the systems include it.
@@ 5.2.2 Time
In contrast, nearly all of the systems include the concept of time. Observation time is the most common way time is incorporated in the systems, with many systems capturing the exact timestamp for each sensor observation [80, 81, 83, 84, 85, 86, 89, 93, 95, 99, 100, 103, 105, 74, 107, 108, 75, 111]. Besides observation time, the time at which certain events occur can be captured, for example calls to emergency services [80]. This allows the systems to display or analyse trends over time. Additionally, Alti et al. [80] capture the time intervals in which reports should be sent. Rather than a timestamp, several systems also capture the general time of day during which observations or activities occur. For instance, Ali et al. [77] divide the time at which daily activities are done into morning, afternoon, and evening. Peral et al. [104] use mealtimes as a point of reference, which is particularly important when taking blood glucose measurements. They distinguish between pre-breakfast, pre-lunch and pre-dinner readings.
Duration and frequency are other important aspects of time. Duration can be captured for physical activity [81, 106], sleep [81, 74, 71], symptoms [101], and treatment [107, 72], and disease [75]. De Brouwer et al. [83] capture headache duration as well as the duration of events that influence headaches, such as stress and sleep. Symptom duration can influence the risk for certain illnesses, while specifying treatment duration ensures medication reminders are sent only during the prescribed period. When combined with thresholds, duration can be useful in identifying different situations. For example, Stavropoulos et al. [74] determine that an individual has a lack of movement if they have fewer than 500 steps and their heart rate has been less than 100 beats per minute for longer than 800 minutes. Frequency is used by Chiang and Liang [82], Spoladore et al. [106], and Yu et al. [110] as a metric for physical activity. Mezghani et al. [102] capture the frequency of sensor observations, while Villarreal et al. [109] capture the frequency of detected diseases.
[…]

### v3

@@ 5.1.5 The role of Semantic Web technologies
Semantic Web technologies do not inherently provide support for technical interoperability, since they operate at a higher, more abstract level to formally represent and derive meaning from the data. Therefore, in order to achieve technical interoperability, health monitoring systems must leverage data transmission protocols and devices. However, Semantic Web technologies are critical in the achievement of semantic interoperability among the selected systems. 20 of the systems make use of terminologies such as SNOMED CT and ICD through ontologies, and access knowledge graphs such as DrugBank and Diseasome which are published as linked data. This allows the systems to access and reason with standardised health domain knowledge.
Semantic Web technologies can also contribute to syntactic interoperability. For instance, El-Sappagh et al. [116] and Kilintzis et al. [97] map FHIR resources to an ontology, allowing for interoperability between their proposed system and hospital information systems that use FHIR. These are the only two systems that use Semantic Web technologies to achieve interoperability with established health standards. This can be attributed to the historical gap in user-friendly tools for this purpose. For example, although FHIR has an RDF representation serialised in Turtle format, early adopters noted several issues precluding its ease of use, such as literal values and FHIR references being nested under blank nodes, and unnecessarily long predicate names [144]. However, more recent versions of FHIR RDF have largely addressed these issues [144].
Finally, with regard to the representation of heterogeneous and multimodal data, 42 systems use ontologies for this purpose. The exceptions are three systems which represent the data in knowledge graphs [108, 105, 104], and three systems which only use Semantic Web technologies (i.e. domain ontologies, external knowledge graphs, and linked data repositories) as a source of domain knowledge [110, 120, 103].
5.2 Situation detection
A situation can be understood as a higher-level interpretation of sensor data that is relevant and of interest in an application domain [4]. Personal health monitoring systems should be capable of situation analysis, which entails both the detection and the prediction of health situations. We discuss situation detection below, and situation prediction in Section 5.3.
5.2.1 Forms of situation detection
In health monitoring systems, situation detection can take a variety of forms. One of these is the categorisation of individual sensor observations based on whether they are within or outside a given range as determined by domain knowledge. For example, in the system proposed by Akhtar et al. [109], when vital signs such as temperature and heart rate are outside the normal range, the situation is classified as an emergency. Likewise, Elhadj et al. [89] classify expected observations as normal, while observations outside the normal ranges are classified as abnormal. They also include a third classification, wrong, for faulty observations from malfunctioning sensors. Similar threshold-based situation categories are used in 19 of the systems [108, 87, 97, 125] [113, 115, 92, 118, 119, 95, 121, 124, 130, 83, 100, 102, 107, 132, 133]. Thresholds have also been used to classify physical activity based on level of intensity [114, 116, 90, 92, 95]. A better approach than using individual sensor observations is to consider different observations and personal attributes to classify individuals. This is done by Ali et al. [112], who classify the patient health condition as either healthy, moderate, or serious based on multiple sensor outputs and properties such as sex, weight, and height. Similarly, Chiang and Liang [88] classify situations as either healthy, moderate, or severe based on age, blood pressure, blood glucose, heart rate, and cholesterol.
Another form of situation detection in health monitoring is the detection of medical conditions and diseases. Some conditions such as hypertension and hyperglycemia can be diagnosed based on individual sensor observation thresholds. This is done by Kim et al. [122], who detect prehypertension and step 1 and 2 hypertension based on defined blood pressure thresholds. Similarly, hyperglycemia is detected by Rhayem et al. [83] based on blood glucose levels. Other diseases require the analysis of signs and symptoms based on a combination of different sensor observations and other sources of data. For example, Ivașcu et al. [96] detect mental disorders (Parkinson’s, Alzheimer’s, psychosis, and depression) using signs and symptoms related to behaviour, motor skills, cognitive skills, facial appearance, mood, sleep, weight, and speech. Other systems are able to detect types of headaches [115], heart disease [111], diabetes [110, 112, 94], frailty [91], stroke [120], and skin and kidney diseases [123]. Beyond the detection of diseases, Zafeiropoulos et al. [107] detect medication adherence in Parkinson’s disease patients by recognising missed doses based on two features: tremors and bradykinisia.
5.2.2 Techniques for situation detection
39 of the systems implement some form of rule-based reasoning for situation detection. Of these, 28 systems specify using Semantic Web rules in particular [109, 112, 113, 87, 114, 88, 116, 89, 90, 92, 118, 93, 94, 119, 97, 122, 124, 128, 129, 106, 83, 98, 99, 100, 131, 107, 132, 133]. A discussion of the specific rule languages used in the systems can be found in Section 6.2.2. Rules provide a way to implement expert knowledge in an if-then form, whereby if certain conditions are met, then a consequent conclusion is made or action taken. Despite their widespread use, rules have several limitations. Firstly, crisp rules are unable to handle uncertainty and ambiguity in sensor observations and the determination of health situations. To mitigate this, five systems incorporate fuzzy logic [112, 88, 90, 117, 129] while one incorporates defeasible logic [109] within the rules. These techniques are discussed in greater detail in Section 5.7, which focuses on techniques for handling uncertainty in health monitoring. Secondly, rules are typically based on existing expert knowledge, and therefore cannot incorporate new knowledge that experts may be unaware of. Additionally, manually updating rules is time-consuming, making them difficult to scale. This challenge can be overcome using learned rules based on ML algorithms, which can acquire new, high quality knowledge automatically [145] and contribute to dynamic and adaptable rule-based systems. The systems proposed by Hussain et al. [120], Henaien et al. [93] and Peral et al. [130] extract rules from decision trees. However, caution should be exercised when using ML-derived rules, as they may still need verification and validation from domain experts. As an alternative to rule-based reasoning, Xu et al. [105] implement case-based reasoning, arguing that it is easier to capture human experiences using cases rather than rules. By searching for historical cases that are similar to the current case, their proposed system is able to obtain treatment plans that have been successful in the past.
In addition to the development of rules as discussed above, ML is also used in a number of the systems for the classification of diseases based not only on sensor data but also other data sources. Ali et al. [110] use a bidirectional long short-term memory (BiLSTM) model to detect diabetes and blood pressure, to classify sentiments from social networking data for mental health monitoring, and to classify drug side effects. Their proposed system uses domain ontologies to extract important features that can enhance the ML classification. Zhou et al. [103] also use a BiLSTM model for disease prediction, while Garcia-Moreno et al. [91] use k-nearest neighbours to classify elderly individuals based on frailty and dependence. Other ML algorithms used include multi-layer perceptron for heart disease detection [111] and random forest for stroke detection [120]. ML is also used for physical activity classification, for example using the k-nearest neighbours [92, 126], decision trees [126], and random forest [95, 126] algorithms. Finally, ML can also be used to classify situation severity for reporting purposes. This is done by Zafeiropoulos et al. [107], who use a graph neural network to distinguish between medium and high alerts. A full review of ML techniques for situation analysis in the health domain is outside the scope of this study. Readers are referred to the reviews by Ravì et al. [146] and Li et al. [147].
5.2.3 The role of Semantic Web technologies
Semantic Web technologies can support situation detection in two main ways: firstly, they formally represent important concepts and the relationships between them, i.e. sensor data, domain knowledge, contextual information, and even the situations themselves; and secondly, they support reasoning through which new knowledge can be derived from existing knowledge [4]. Although several situation-focused ontologies have been developed, including the Situation Theory Ontology [148] and the Scenes and Situations ontology [149], none of the selected systems reuse any such ontologies. Despite this, Semantic Web technologies remain vital for situation detection among the selected systems. Rule-based reasoning is the most common mechanism for situation detection among the systems. These rules rely heavily on concepts that are formally defined in ontologies, and they are more often than not expressed in standard Semantic Web languages such as SWRL.
5.3 Situation prediction
All 48 selected systems detect current situations. In contrast, only 12 of the systems go beyond this to predict some future outcome [111, 113, 88, 115, 117, 119, 127, 130, 106, 83, 131, 103].
5.3.1 Forms of situation prediction
All 12 of these systems explore the concept of risk as a situation prediction feature, since determining an individual’s risk profile for a certain condition can be used to predict future adverse health situations. This includes the risks of heart disease [110, 119], arthritis recurrence [88], stroke [127], and fetal loss in gestational diabetes patients [83]. Zhou et al. [103] use a multi-label classification approach to simultaneously assess the risk of multiple chronic diseases, including hypertension, diabetes, and arthritis. De Brouwer et al. [115] detect triggers as a means to anticipate potential headache attacks in the future, which can be considered a form of risk prediction. To support the identification of potential risks, future physiological readings can also be predicted using historical sensor observations, as is done by Peral et al. [130]. Their proposed system predicts blood glucose levels over three-day and five-day windows. These predictions of sensor measurements can then be analysed to determine future health risks. Hristoskova et al. [119] similarly adopt a temporal analysis, determining the risk of congestive heart failure over a four-year time horizon.
Another useful aspect of situation prediction is the determination of the prognosis, i.e. expected progression, of a detected disease, although this is poorly explored in the systems. Hussain and Park [120] mention the intention to extend their system in future work to include automated stroke prognosis; however, this is not implemented in the current version of the system. In contrast, Yu et al. [131] include a disease progression class in their proposed ontology, representing past diagnoses or potential health risks and their associated times. However, the system does not include any methods to predict the progression of detected conditions.
5.3.2 Techniques for situation prediction
Similar to situation detection, the most common technique used in situation prediction is rules, which is used in nine of the 12 systems. Of these, seven specify Semantic Web rules in particular [113, 88, 119, 106, 83, 131, 107]. As discussed in Section 5.2, both Chiang and Liang [88] and Fenza et al. [150] use fuzzy rules to enhance crisp rules. ML is used for situation prediction in 4 of the systems [111, 130, 107, 103]. Ali et al. [111] use an ensemble deep learning classifier, which consists of a five-layer feed-forward network that incorporates a boosting algorithm, to predict future heart attacks, while Peral et al. [130] use support vector machine and logistic regression models to forecast future blood glucose measurements. Although De Brouwer et al. [115] use ML to detect headache triggers, the actual prediction of headaches is knowledge-based using SPARQL queries on stored data.
Besides rules and ML, Bayesian networks are used in two of the systems. These are probabilistic models in the form of directed acyclic graphs that can represent causal relationships among variables in a domain. Mcheik et al. [127] use a Bayesian network to calculate the risk of stroke occurring in the next seven days, based on risk factors such as age, presence of diabetes, high blood pressure, and symptom duration. This approach is also taken by Kordestani et al. [123] to determine the probability of the occurrence of kidney disease. The use of fuzzy rules and Bayesian networks is discussed in greater detail in Sections 5.6 on explainability and  5.7 on uncertainty handling.
5.3.3 The role of Semantic Web technologies
By providing a structured representation of risk factors, temporal concepts, health outcomes, and situations, and supporting reasoning over these concepts, Semantic Web technologies support more effective situation prediction. Similarly to situation detection, rules remain the most used reasoning approach for situation prediction among the systems, with seven of them using Semantic Web rules specifically. However, as discussed in Section 5.2, semantic-based approaches to situation analysis face drawbacks such as scaling difficulties and the inability to handle uncertainty inherently. These limitations can be mitigated by combining them with complementary techniques such as ML, fuzzy logic, and Bayesian networks. In fact, several extensions of Semantic Web standards and languages have been proposed that incorporate these uncertainty handling techniques, and these are discussed in Section 5.7.6. However, none of these are used in the reviewed systems.
5.4 Decision Support
Decision support is the natural next step after situation analysis. Based on the detected and predicted situations, targeted support can be offered to mitigate adverse situations and promote favourable health outcomes. 45 systems implement some form of decision support.
5.4.1 Forms of decision support
Messages and notifications are used in majority of the systems to warn of potentially dangerous situations, issue reminders, and prompt mitigating action. These notifications can be sent to monitored individuals, caregivers, clinicians, and emergency services depending on the severity of the situation. The most common form of decision support is alerting users of adverse situations. This is done in nearly all of the selected systems, with the exception of 11 which do not mention this crucial functionality [110, 111, 113, 114, 116, 117, 97, 122, 129, 106, 131]. A well-documented issue with alerts in the health domain is alert fatigue, a phenomenon in which users become desensitized to alerts due to their frequency [151]. Esposito et al. [90] mitigate this by differentiating between critical and non-critical abnormal situations, with the latter being sent out in a daily summary report email rather than an instantaneous notification for each case. Additionally, 10 systems send reminder messages of some kind to users, for instance about medications, exercise, and other interventions [108, 88, 118, 93, 123, 126, 99, 100, 101, 104]. For example, the system proposed by Kordestani et al. [123] can remind clinicians to order additional laboratory tests when previously taken tests become out of date.
In addition to alerts and reminders, health monitoring systems can also trigger actions in response to adverse situations. For example, the system proposed by Alti et al. [113] triggers the injection of insulin in response to a blood glucose level above a certain threshold, while the system proposed by Hadjadj and Halimi [118] can trigger the opening of a vehicle door. Such systems must be integrated with an actuation device capable of carrying out the action. The system proposed by Titi et al. [100] includes several actuators such as a smoke alarm. Other systems are integrated with actuators capable of opening doors and moving beds [87], turning on lights and heaters [88], making emergency calls [83], or turning off water or gas if detected [101].
Another form of decision support is the generation of suggestions or recommendations for the mitigation of adverse situations, which is implemented in 29 systems. The recommendations include lifestyle modifications, such as diet and activity [111, 112, 113, 108, 114, 116, 92, 118, 120, 122, 124, 83, 98, 102, 104], medication [113, 116, 89, 119, 123, 124, 130, 83, 100, 105], suitable environmental conditions [88], hospital visits [103], and other unspecified treatments and mitigations [110, 109, 115, 125, 126, 107]. Two important considerations when choosing appropriate medications are their side effects and how they interact with other medications. Ali et al. [110] use drug review websites to collect data on side effects, while Elhadj et al. [89] keep track of medication interactions as well as patient allergies. Similarly, the system proposed by Ammar et al. [108] includes an app where monitored individuals can report medication side effects. This information assists clinicians in prescribing appropriate medications for each patient. Related to recommendations is the ability for the monitored individual to seek out relevant and trusted medical information. For example, the system proposed by Rhayem et al. [83] includes a notification module that allows patients to contact a clinician and receive recommendations and treatments from them.
5.4.2 Techniques for decision support
Much like situation detection and situation prediction, rule-based reasoning is the most commonly used technique for decision support in the selected systems. Rules are used by 37 of the 45 systems that implement decision support, of which 27 systems use Semantic Web rules in particular [109, 111, 112, 113, 87, 114, 88, 116, 89, 90, 92, 118, 93, 94, 119, 97, 122, 124, 128, 83, 98, 99, 100, 131, 107, 132, 133]. They are typically implemented based on detected and predicted situations, such that adverse situations trigger alerts or recommendations. Zafeiropoulos et al. [107] combine ML and rules, using a graph convolution network to classify alerts into one of two levels (medium or high), after which the alerts are sent based on rules.
Similar to their approach for situation detection, Xu et al. [105] adopt case-based reasoning rather than rules for decision support. They compare treatments and clinical paths in similar patients, and, if differences are found, the system sends an alert to flag the treatment plan for adjustment. Among the agent-based systems, the functionality related to decision support is often delegated to agents, such as the notification and alert agents in the systems proposed by Ivașcu et al. [96] and Ivașcu and Negru [95]. Additionally, interactive agents and chatbots play an important role in decision support. Ammar et al. [108] propose digital assistants that provide personalised alerts and suggestions and summarise text. Similarly, Mavropoulos et al. [126] propose a smart virtual agent that clinicians can interact with via voice commands, while Yu et al. [104] implement an AI chatbot to answer user questions.
5.4.3 Quality of decision support
Furthermore, the soundness of recommended interventions can be ensured by incorporating established and clinically validated medical guidelines or vetted information from medical or allied health professional bodies. This can contribute to the acceptance of health monitoring systems by medical professionals and regulatory bodies. Despite this, only 15 of the selected systems mention the use of such guidelines or vetted medical information. They include Lopes de Souza et al. [124], who incorporate risk level classifications from the American Heart Association, and De Brouwer et al. [115], who use the International Classification of Headache Disorders’ criteria to issue relevant alerts. Garcia-Moreno et al. [91] and Spoladore et al. [98] use the International Classification of Functioning, Disability and Health for health status classification; while Ali et al. [111] and Hristoskova et al. [119] use the Framingham Risk Score to determine congestive heart failure risk. El-Sappagh et al. [116] extract knowledge from bodies such as the American and Canadian diabetes associations, while Yu et al. [131] use guidelines from the United Kingdom’s National Health Service. Finally, the system proposed by Ammar et al. [108] includes an option for clinicians to refer patients to trusted medical knowledge sources such as the Centers for Disease Control and Prevention (CDC).
5.4.4 The role of Semantic Web technologies
The main value that Semantic Web technologies add to decision support in the selected systems is their ability to represent important concepts, particularly situations and domain knowledge, which can then be reasoned over. This is consistent with our findings on the role of Semantic Web technologies in situation detection and situation prediction. It also aligns with previous research findings on the use of Semantic Web technologies for decision support [13] [154]. Rules are the most common reasoning tool among the selected systems, with Semantic Web rules used for decision support in 27 of them. However, rule-based reasoning for decision support faces the same limitations in scalability and lack of inherent uncertainty handling as discussed in the previous sections.
@@ 5.5 Context awareness
An important aspect of health monitoring is the ability to take context into consideration, which is critical for situation analysis since contextual information enhances sensor data and supports its interpretation. This, in turn, impacts decision support. Consider a case where an individual’s heart rate is suddenly elevated. If the individual is engaged in exercise, the increased heart rate is expected. However, if the individual is at rest, this could be a cause for alarm and may necessitate intervention. Therefore, health monitoring systems must be able to adapt based on the context of the individual being monitored. The four most common aspects of context are location, time, identity (of a person or agent), and activity (or events) [155, 156].
@@ 5.5.1 Location
Ye et al. [155] highlight three types of locations that can be represented: symbolic locations, coordinate locations, and regions. The systems proposed by Akhtar et al. [109], Bampi et al. [87], Chiang and Liang [88], and Vadillo et al. [101] keep track of the different rooms in a house where an individual may be, while those proposed by Khozouie et al. [121], Titi et al. [100], and Zeshan et al. [132] indicate more generally the place the monitored individual is (for example, “home” or “hospital”). These are symbolic locations. One purpose of such locations is to allow the systems to suggest relevant services based on the type of space currently occupied, as is the case in the system proposed by Chiang and Liang [88]. In the system proposed by Hristoskova et al. [119], the clinician’s location (i.e. the room they occupy in a hospital) is used to determine which device to send notifications to, optimising for the closest device. Likewise, Zeshan et al. [132] determine the closeness between the monitored individual and their caregivers in order to select which caregiver or clinician to notify in case of abnormal sensor observations. This is similar to the system proposed by Alti et al. [113], which supports a GPS sensor that captures the current coordinates of the monitored individual. In this system, location is used to select devices closest to the user from which to deploy health services so as to increase efficiency and minimise inter-device communication costs. Coordinate locations also serve the purpose of alerting caregivers and emergency services of the exact location of a person in the event of a medical emergency, as is suggested by Rhayem et al. [83]. The system proposed by Hadjadj and Halimi [118] integrates health monitoring in the public transport system, and therefore includes location sensors in public transportation vehicles. The final type of location is regions, which are geometrical two- or three-dimensional representations of locations [155]. This type of location is used in the system proposed by Kim et al. [122] in order to advise users of region-specific situations, such as adverse or dangerous weather. Similarly, Ammar et al. [108] use ZIP codes, which represent a small geographic region, to extract neighbourhood-specific social determinants of health such as walkability. El-Sappagh et al. [116] use the spatial region class from the Basic Formal Ontology to represent the patient’s current location, as well as the placement of the sensors. Despite the importance of location as an aspect of context, only 21 of the 48 systems [109, 113, 108, 87, 88, 115, 116, 89, 91, 118, 93, 119, 121, 122, 125, 106, 83, 100, 101, 131, 132] include it.
@@ 5.5.2 Time
In contrast, all of the systems include the concept of time with the exception of five [111, 112, 117, 94, 104]. Observation time is the most common way time is incorporated in the systems, with 25 systems capturing the exact timestamp for each sensor observation [87, 91, 97, 125] [113, 114, 115, 116, 89, 90, 118, 120, 96, 121, 124, 126, 129, 130, 106, 83, 99, 100, 101, 107, 132]. Besides observation time, the time at which certain events occur can be captured, for example calls to emergency services [113]. This allows the systems to display or analyse trends over time. Additionally, Alti et al. [113] capture the time intervals in which reports should be sent. Rather than a timestamp, Ali et al. [110] record the general time of day during which daily activities occur, i.e. morning, afternoon, or evening. Similarly, Peral et al. [130] use mealtimes as a point of reference, which is particularly important when taking blood glucose measurements. They distinguish between pre-breakfast, pre-lunch and pre-dinner readings.
Duration and frequency are other important aspects of time. Duration can be captured for physical activity [114, 98], sleep [114, 99, 103], disease [107], symptoms [127], and treatment [100, 105]. De Brouwer et al. [115] capture headache duration as well as the duration of events that influence headaches, such as stress and sleep. Symptom duration can influence the risk for certain illnesses, while specifying treatment duration ensures medication reminders are sent only during the prescribed period. When combined with thresholds, duration can be useful in identifying different situations. For example, Stavropoulos et al. [99] determine that an individual has a lack of movement if they have fewer than 500 steps and their heart rate has been less than 100 beats per minute for longer than 800 minutes. Frequency is used by Chiang and Liang [88], Spoladore et al. [98], and Yu et al. [131] as a metric for physical activity. Other examples of frequency in the systems are frequency of sensor observations [128] and frequency of disease occurrence [102].
Notably, valuable features can be extracted from changes in time series sensor data. For instance, Hussain and Park [120] and Ivașcu and Negru [95] use the time-domain features of the ECG to calculate heart rate and heart rate variability. Additionally, the multi-agent system proposed by Akhtar et al. [109] incorporates temporal logic, which allows for the formalization of temporal ordering operators such as “next”, “always”, “until”, and “while” without referencing actual times [157]. Another interesting time-related aspect is trajectory, which combines both spatial and temporal properties to represent the mobility of a sensor. This is incorporated in the system proposed by Rhayem et al. [83] to define a source and destination of a sensor within a particular duration of time.